OpenStudy (anonymous):

A weight lifter struggles but manages to keep a heavy barbell abovehis head. Occasionally he slips and the barbell starts to fall downward, but he always recovers. In each case, compare the force exerted by the weight lifter on the barbell to that exerted by the barbell on the weight lifter. A)- With the barbell at rest. B)- With the barbel moving upward. C)- With the barbel moving downward.

OpenStudy (amriju):

equal...action reac forces

OpenStudy (amriju):

and in all cases

OpenStudy (anonymous):

can you explains that @amriju ??

OpenStudy (amriju):

Frce exerted by the man on barbell will be responded by an equal and opposite force on the man by thebarbell....newtons 3rd

OpenStudy (amriju):

thats irrespective of what the case may be...

OpenStudy (amriju):

when this force exceeds the force of gravity on the barbl...it gets risen up

OpenStudy (anonymous):

The reaction force in each case equals to the weight of the barbell. The difference is in the lifting force by the arms of the weightlifter. A)Net force=0. So they are the same. B)Net force=upwards force: Weight of barbell<force by weight lifter.(so that he can lift it) C)Net force=downwards. Weight of barbell>force by weight lifter.(he can't lift it)

OpenStudy (amriju):

how can u say that the reaction force equals weight of barbell?...it equals the action...thats the force of the weight lifter on the barbell

OpenStudy (anonymous):

The reaction force considered is the reaction force on the barbell. Since The force exerted by the barbell=it's weight, W The force on the barbell=-W+F+R(=W) The force exerted by the weight lifter = F The force on the hands=W+R(=F)

OpenStudy (anonymous):

The barbell only gives out one force by it self and that is its weight